OpenStudy (anonymous):

solve the quadratic inequality. write the solution set in interval notation. x^2- 9x+ 18> 0

OpenStudy (anonymous):

i got \[-1\pm 9\sqrt{1}\] is that correct

OpenStudy (anonymous):

can someone please help me

OpenStudy (anonymous):

You need to use the Quadratic formula which is, http://math.about.com/od/Equations/ss/Quadratic-Function-6_2.htm

OpenStudy (anonymous):

i know that but im not sure if i did it correct

OpenStudy (anonymous):

show me what you got

OpenStudy (anonymous):

\[−1±9\sqrt{1}\]

OpenStudy (anonymous):

I got 9 and 6 \[-(-9)\pm \sqrt{(-9)^2-4(1)(18)} all \over (2)(1) \]

OpenStudy (anonymous):

how did you get 9 and 6?

OpenStudy (anonymous):

i got 6 and 3

OpenStudy (anonymous):

why is the 3 -

OpenStudy (anonymous):

because i used subtraction. the (-) just to show me which sign i have used.

OpenStudy (anonymous):

Btw the answer is (+)

OpenStudy (anonymous):

oh okay thank you sooo much

OpenStudy (anonymous):

no Problem and good luck